# On the polemic assessment of what Bell did

### v2

@@
On the polemic assessment of what Bell did
Justo Pastor Lambare jupalam@gmail.com Universidad Nacional de Asunción-FACEN, Ruta Mcal. J. F. Estigarribia, Km 11 Campus de la UNA, San Lorenzo-Paraguay
(9 de septiembre de 2025)
Resumen
Despite their Nobel Prize-winning empirical falsification, the interpretation of the Bell inequality remains a subject of controversy. This article discusses and attempts to clarify the reasons John S. Bell and A. Einstein claimed that quantum entanglement implies puzzling nonlocal correlations that Einstein famously termed “spooky action at a distance.” The issue remains highly controversial and has roughly divided the scientific community into localists and nonlocalists. Without taking a stance for either side in the long-standing, polarized debate, we examine Bell’s actual argument, highlighting the differences between his approach and the current orthodoxy.
I Introduction
Bell, of course, like Einstein, believed that quantum mechanics implies “Spooky action at a distance.” However, contrary to a widespread opinion, we sustain that Bell did not claim that the proof of quantum nonlocality should involve his “locality inequality.”
The intention of this work is not to challenge the current dominant interpretations on the locality problem, but to expose Bell’s true ideas by analyzing his writings with an objective and undogmatic attitude, disentangling what Bell’s argument indeed was and highlighting the subtle yet important conceptual differences with the prevailing orthodoxy.
@@ I Introduction
The scientific community’s interpretation is generally divided into localists and nonlocalists; however, both groups invariably base their arguments on the precept that the proof of quantum nonlocality relies on the violation of the Bell inequality. A characteristic exposition of this dichotomy is well depicted by an exchange between Tim Maudlin, with his article “What Bell did” which suggested the title of this paper, and Reinhard Werner’s comments, exemplifying the innumerable articles claiming quantum nonlocality based on the Bell inequality, on the one hand, and its rejection for the same reason, on the other [1, 2, 3, 4].
In section II we present evidence supporting that Bell did not claim his inequality is proof of quantum nonlocality. The section has a historical character introducing a chronological review of Bell’s main papers on nonlocality. It shows that, in his most lucid writings, Bell distinctly formulated his inequality only after explicitly establishing the nonlocal character of quantum mechanics, either by previously assuming [5] or directly proving it [6, 7].
Section III delves into Bell’s actual quantum nonlocality argument, which is based on the local causality concept appearing in 1975, more than ten years after the publication of his celebrated 1964 paper. We highlight a formal application of local causality to quantum mechanics (sec. III.1.2), and identify the origin of the longstanding dispute between localists and nonlocalists (sec. III.1.3).
@@
    The $\lambda$ variables, understood as common causes, are not necessarily “elements of physical reality” in the EPR sense [8]. Indeed, although Bell often referenced the EPR paradox, he never resorted to the concept of elements of physical reality. Moreover, he explicitly mentioned that the $\lambda$ variables could well include the quantum state vectors (cf. sec. III.1.2 and appendix B).
    Thus, according to Bell’s reasoning, the $\lambda$ variables are fairly general parameters. They are not conditioned to be preexisting entities, are wholly compatible with quantum “nonrealism” [^1] By realism in this paper, we mean the usual connotation given in physics, i.e., the view according to which physical magnitudes are assumed to exist and have definitive values whether or not they are observed. and do not necessarily disrupt quantum superposition.
@@
Bell’s local causality concept aligns with the Principle of Common Causes, which was independently and more completely developed by Reichenbach [9] to study the problem of causal structures and probabilistic correlations. According to Cavalcanti and Lal [10]:
@@
II Chronological Review of Bell’s Works on Nonlocality
@@
On the other hand, he explicitly proved quantum mechanics violates local causality before formulating any inequality at least on two occasions [6, 7]. Therefore, contrary to the usual folklore, the objective textual evidence indicates that Bell did not consider his inequality as a direct proof of quantum nonlocality but only of the impossibility of a non-conspiratorial local completion. [^2] Non-conspiratorial means that the statistical independence condition is assumed (cf. III.2).
II.1 The 1964 Bell Theorem
According to the prevalent view, Bell formulated his inequality to prove that quantum mechanics is not a local theory, presenting the Bell theorem as a quantum nonlocality theorem. But an attentive reading of Bell’s 1964 argument reveals two crucial facts that are widely overlooked and prove otherwise [5]:
@@ II.1 The 1964 Bell Theorem
Bell resumed the argument where EPR has left it: if local, quantum mechanics must be incomplete. Since the orthodox interpretation asserts quantum mechanics is complete, then, according to EPR, it must be nonlocal (cf. sec. III.3).
The previous inference was implicit in his expression referenced in a), asserting that additional variables were necessary to restore locality. Then, applying an EPR-like reasoning, Bell derived a deterministic hidden variable model and proved, through his inequality, the untenability of a local completion. [^3] Recently, Michael Hall claimed that deterministic hidden variables cannot be derived from Bell’s 1964 hypotheses. However, as Hall himself recognizes, the issue is not relevant because later generalizations of the inequality did not assume perfect correlations; besides, stochastic Bell inequalities were formulated [11].
@@
Bell’s theorem is a simple mathematical theorem that should be free of any polemic if we strictly follow Bell’s rationale, namely, that a local completion is untenable. Probably, that was what Richard Feynman meant when he said that Bell’s theorem [12],
@@
However, we disagree with Feynman on the unimportance of the Bell theorem since it was a significant advance in the Bohr-Einstein debate that, by 1964, remained stagnant for almost thirty years.
II.2 Bell’s theorem after 1964
Bell’s arguments evolved over the years. In later papers, although he occasionally mentioned EPR, he mostly abandoned the EPR reasoning, and whenever explicitly proving quantum nonlocality, he always used the concept of local causality. As in his 1964 paper, in the articles he published afterward, we could not find any convincing evidence suggesting that he considered his inequality to be a proof of quantum nonlocality. However, there is clear evidence to the contrary.
Although he generalized the “locality inequality” to a stochastic version, there is sufficient evidence indicating how he continued interpreting his inequality as proving only the impossibility of a local completion. Indeed, at least on two occasions [6, 7], he unambiguously concluded that quantum mechanics is not locally causal before introducing hidden variables and formulating any inequality.
@@ II.2 Bell’s theorem after 1964
II.2.1 Introduction to the hidden variable problem
In 1971, he wrote the paper “Introduction to the hidden-variable question” [13]. Here, Bell did not address the question of quantum nonlocality but, as the title explicitly reveals, investigated hidden variable theories using the de Broglie-Bohm theory as a paradigmatic example, stressing its conspicuous nonlocal character as the “difficulty.”
His paper highlights the fact that de Broglie-Bohm theory is not nonlocal incidentally, but that any extension of quantum mechanics must necessarily be nonlocal. Bell proved that quantum mechanics violates his inequality, concluding:
@@ II.2.1 Introduction to the hidden variable problem
Note that the adjective “local” in his expression refers to the hidden-variable theory, so irrespective of whether he believed that quantum mechanics is nonlocal, he did not say his inequality proved it. Any assertion in that direction would be unfounded and contrary to what he indeed used to explain elsewhere when explicitly proving quantum nonlocality.
II.2.2 The theory of local beables
This article appeared in 1975.444Bell’s work is reproduced in [6]. Here, Bell abandoned the EPR reasoning and set out to prove quantum nonlocality explicitly by introducing the concept of local causality. He argued that quantum mechanics violates this form of locality in section 3 without mentioning any inequality. He starts that section by asserting:
@@ II.2.2 The theory of local beables
(2) above refers to the local causality constraint, a concept he developed in the previous section. Then, he proceeded to present his argument on quantum nonlocality. In the same section, but only after establishing the nonlocal character of quantum mechanics, Bell explored the problem of adding hidden variables.
Then in section 4, “Locality inequality,” he derived a stochastic Bell-CHSH inequality. Finally, in section 5, he established the impossibility of a local completion by proving that quantum mechanics violates his inequality, concluding:
@@
That is different from concluding, “So quantum mechanics is not a local theory” unless we force the interpretation by ignoring his procedure. Otherwise, why would he bother to first prove quantum mechanics violates local causality two sections before without using any inequality or hidden variables? We guess because he agreed with Stapp about the logical loophole of concluding quantum nonlocality directly from his inequality [14]:
@@
Above, “properties alien to the original system” rigorously mean variables that do not legitimately pertain to quantum mechanics. Although some have observed that the hidden variables can include the quantum state [15, 16, 17], the problem persists with the other “additional variables.” As we observe in the section III.1.3, the Bell inequality cannot be formulated without additional variables foreign to quantum mechanics, notwithstanding that one of those variables may include the quantum state as indicated by Bell himself [18].
II.2.3 Bertlmann’s socks
In 1981 Bell wrote his celebrated paper “Bertlmann’s socks and the nature of reality” [18]. On this occasion, Bell did not explicitly prove quantum nonlocality before formulating the inequality. He based his arguments on EPR. But he also used common causes to explain local correlations.
This is one of the few papers where there is some ambiguity on whether he considered his inequality violation as proof of quantum nonlocality. Can we assume that Bell changed his mind about the meaning of his inequality? We do not think so because, in his last paper (cf. II.2.4), he returned to his previous formulations, i.e., either assuming [5] or proving [6] quantum nonlocality without introducing hidden variables or mentioning any inequality.
In “Bertlmann’ socks”, Bell chose intuition and ease of interpretation over logical rigor. In Bell’s own words, this paper was one of those that [19]:
@@ II.2.3 Bertlmann’s socks
That is why he spent great effort explaining the difference between quantum and classical entanglement through naive analogies, such as those of the socks of Mr. Bertlmann. It is rather surprising that some scientists still believe that quantum entanglement admits such a trivial explanation (cf. sec. A.2).
II.2.4 La nouvelle cuisine
This is Bell’s last paper which appeared in 1990 [7]. Here again, Bell’s view of his inequality and quantum nonlocality is crystal clear.
This time his approach differs from his 1975 formulation in “The theory of local beables” in that he proves that quantum mechanics violates local causality through the EPR reasoning (cf. III.3). However, what remains unchanged is that he proved quantum nonlocality in section 8 before formulating his inequality.
@@ II.2.4 La nouvelle cuisine
Although some interpret the above expression as proving that nature itself, including quantum mechanics, is nonlocal, the order in which he presents his argument is unambiguous. First, establishing quantum nonlocality without any inequality, and then, only afterward, in a separate section, through his inequality, inferring his conclusion about extended locally causal theories.
@@
II.2.5 Further writings
In 1975, an editor asked Bell to write a response letter [20]. He started the paper by expressing:
@@ II.2.5 Further writings
In 1977 Bell wrote Free variables and local causality [21]. He started the paper by writing:
@@
Once more, his expression openly shows his reasoning. He considered the statement “quantum mechanics is not locally causal” to be before, different from, and not a consequence of, the assertion “cannot be embedded in a locally causal theory” which was the conclusion he used to draw from the violation of his inequality as an additional characteristic of quantum mechanics besides nonlocality, as he explicitly did it in Refs. [6, 7] and we explain in sections II.2.2 and II.2.4.
III Bell’s Proof of Quantum Nonlocality
In this section, we analyze Bell’s explicit argument on the nonlocal character of quantum mechanics. In his 1975 paper, “The theory of local beables” [6], Bell gave an explicit argument for quantum nonlocality for the first time. This paper has four outstanding characteristics that were missing in 1964:
@@ III Bell’s Proof of Quantum Nonlocality
III.1 Local Causality and Quantum Nonlocality
Firstly, we briefly review the meaning of LC. Unfortunately, as observed by Norsen [15], LC is a little-known concept, and we would add that, if known, it is often misunderstood. For this reason, appendix A contains a more detailed explanation for those unfamiliar with it. Secondly, like Bell, we prove that quantum mechanics violates LC without the Bell inequality. Thirdly, we analyze the origin of the argumentative conflict between localists and nonlocalists.
III.1.1 Local causality
@@
(2) means that after all factors (known and unknown) are included, whatever Bob decides to do in his distant laboratory cannot influence Alice’s local measurements; and vice versa. Therefore, (2) must be satisfied by locally causal theories. Note that $\lambda$ represents local common causes in general, not necessarily “preexisting values” and can well include the quantum state (cf. appendix B).
III.1.2 Quantum nonlocality

### v3

@@
Abstract
Despite their Nobel Prize-winning empirical falsification, the interpretation of Bell’s inequality remains a subject of controversy. This article discusses and attempts to clarify the reasons John S. Bell and A. Einstein claimed that quantum entanglement implies puzzling nonlocal correlations that Einstein famously termed “spooky action at a distance.” The issue remains notoriously controversial and has roughly divided the scientific community into localists and nonlocalists. Without taking a stance on either side in the long-standing, polarized debate, we examine Bell’s actual argument and highlight how his reasoning differs from the current orthodoxy.
Keywords: nonlocality, local causality, Bell’s inequality, common cause principle, statistical independence.
1 Introduction
The intention of this work is neither to challenge the current dominant interpretations nor to solve the long-standing controversy on the nonlocality problem, but to expose Bell’s true ideas by analyzing his writings with an objective and undogmatic attitude, disentangling what Bell’s argument indeed was and highlighting the subtle yet important conceptual differences with the prevailing orthodoxy.
@@ 1 Introduction
Regarding quantum mechanics, the scientific community’s interpretation is generally divided into localists and nonlocalists. However, unlike Bell’s approach, both groups invariably base their arguments on the precept that the proof of quantum nonlocality relies on the violation of the Bell inequality.
A characteristic exposition of this dichotomy is well depicted by an exchange between Tim Maudlin, with his article “What Bell did” which suggested the title of this paper, and Reinhard Werner’s comments, exemplifying the innumerable articles claiming quantum nonlocality based on the Bell inequality, on the one hand, and its rejection for the same reason, on the other [Maudlin 2014b, Werner 2014a, Maudlin 2014a, Werner 2014b].
In section 2 we present evidence supporting that Bell did not claim his inequality is a proof of quantum nonlocality. The section has a historical character introducing a chronological review of Bell’s main papers on nonlocality. It shows that, in his most lucid writings, Bell formulated his inequality only after explicitly establishing the nonlocal character of quantum mechanics, either by previously assuming it [Bell 1964] or by directly proving it [Bell, Shimony, Horne, and Clauser 1985, Bell 2001].
Section 3 delves into Bell’s actual quantum nonlocality argument, which is not based on EPR but on local causality, a concept he introduced in 1975, more than ten years after the publication of his celebrated 1964 paper. We highlight a formal application of local causality to quantum mechanics (sec. 3.1.2), and identify the origin of the longstanding dispute between localists and nonlocalists (sec. 3.1.3).
@@
  • The $\lambda$ variables, understood as common causes, are not necessarily “elements of physical reality” in the EPR sense [Einstein, Podolsky, and Rosen 1935]. Indeed, although Bell often referenced the EPR paradox, he never resorted to the concept of elements of physical reality. Moreover, he explicitly mentioned that the $\lambda$ variables could well include the quantum state vectors (cf. sec. 3.1.2 and appendix 5.3). Thus, according to Bell’s reasoning, the $\lambda$ common causes variables are fairly general parameters. They are wholly compatible with quantum nonrealism, [^2] By realism in this paper, we mean the usual connotation given in physics, i.e., the view according to which physical magnitudes are assumed to exist and have definitive values whether or not they are observed. do not necessarily disrupt quantum superposition, and are not conditioned to be hypothetical values that preexist before actual measurements.
@@
Bell’s local causality concept aligns with the Principle of Common Causes, which was independently and more completely developed by Reichenbach [Reichenbach 1956] to study the problem of causal structures and probabilistic correlations. According to Cavalcanti and Lal [Cavalcanti and Lal 2014]:
@@
2 Chronological Review of Bell’s Works on Nonlocality
@@
On the other hand, he explicitly proved quantum mechanics violates local causality before formulating any inequality at least on two occasions [Bell, Shimony, Horne, and Clauser 1985, Bell 2001]. Therefore, contrary to the usual folklore, the objective textual evidence indicates that Bell did not consider his inequality as a direct proof of quantum nonlocality but only of the impossibility of a non-conspiratorial local completion. [^3] Non-conspiratorial means that the statistical independence condition is assumed (cf. 3.2).
2.1 The 1964 Bell Theorem
According to the prevalent view, Bell formulated his inequality to prove that quantum mechanics is not a local theory, presenting the Bell theorem as a quantum nonlocality theorem. But an attentive reading of Bell’s 1964 argument reveals two crucial facts that are widely overlooked and prove otherwise [Bell 1964]:
@@ 2.1 The 1964 Bell Theorem
Bell resumed the argument where EPR has left it: if local, quantum mechanics must be incomplete. Since the orthodox interpretation asserts quantum mechanics is complete, then, according to EPR, it must be nonlocal (cf. sec. 3.3).
The previous inference was implicit in his expression referenced in a), asserting that additional variables were necessary to restore locality. Then, applying an EPR-like reasoning, Bell derived a deterministic hidden variable model and proved, through his inequality, the untenability of a local completion. [^4] Recently, Michael Hall claimed that deterministic hidden variables cannot be derived from Bell’s 1964 hypotheses. However, as Hall himself recognizes, the issue is not relevant because later generalizations of the inequality did not assume perfect correlations; besides, stochastic Bell inequalities were formulated [Hall 2024].
@@
Bell’s theorem is a simple mathematical theorem that should be free of any polemic if we strictly follow Bell’s rationale, namely, that a local completion is untenable. Probably, that was what Richard Feynman meant when he said that Bell’s theorem [Whitaker 2016],
@@
However, we disagree with Feynman on the unimportance of the Bell theorem since it was a significant advance in the Bohr–Einstein debate that, by 1964, remained stagnant for almost thirty years.
2.2 Bell’s theorem after 1964
Bell’s arguments evolved over the years. In later papers, although he occasionally mentioned EPR, he mostly abandoned the EPR reasoning. Whenever explicitly proving quantum nonlocality, he always used the concept of local causality, which he conceived to be independent of EPR and determinism.
As in his 1964 paper, in the articles he published afterward, we could not find any convincing evidence suggesting that he considered his inequality to be a proof of quantum nonlocality. However, there is clear evidence to the contrary.
Although he generalized the “locality inequality” to a stochastic version, there is sufficient evidence indicating how he continued interpreting his inequality as proving only the impossibility of a local completion. Indeed, at least on two occasions [Bell, Shimony, Horne, and Clauser 1985, Bell 2001], he unambiguously concluded that quantum mechanics is not locally causal before introducing hidden variables and formulating any inequality.
@@ 2.2 Bell’s theorem after 1964
2.2.1 Introduction to the hidden variable question
In 1971, he wrote the paper “Introduction to the hidden-variable question” [Bell 2004]. Here, Bell did not address the question of quantum nonlocality but, as the title explicitly reveals, investigated hidden variable theories using the de Broglie–Bohm theory as a paradigmatic example, stressing its conspicuous nonlocal character as the “difficulty.”
His paper highlights the fact that de Broglie–Bohm theory is not nonlocal incidentally, but that any extension of quantum mechanics must necessarily be nonlocal. Bell proved that quantum mechanics violates his inequality, concluding:
@@ 2.2.1 Introduction to the hidden variable question
Note that the adjective “local” in his expression refers to the hidden-variable theory, so irrespective of whether he believed that quantum mechanics is nonlocal, he did not say his inequality proved it. Any assertion in that direction would be unjustified and contrary to what he indeed used to explain elsewhere when explicitly proving quantum nonlocality.
2.2.2 The theory of local beables
This article appeared in 1975.555Bell’s work is reproduced in [Bell, Shimony, Horne, and Clauser 1985]. Here, Bell abandoned the EPR reasoning and set out to prove quantum nonlocality explicitly by introducing the concept of local causality. He argued that quantum mechanics violates this form of locality in section 3 without mentioning any inequality. He starts that section by asserting:
@@ 2.2.2 The theory of local beables
(2) above refers to the local causality constraint, a concept he developed in the previous section. Then, he presented his argument on quantum nonlocality. In the same section, but only after establishing the nonlocal character of quantum mechanics, Bell explored the problem of adding hidden variables.
Then in section 4, “Locality inequality,” he derived a stochastic Bell–CHSH inequality. Finally, in section 5, he established the impossibility of a local completion by proving that quantum mechanics violates his inequality, concluding:
@@
That is different from concluding, “So quantum mechanics is not a local theory” unless we force the interpretation by dismissing his procedure. Otherwise, why would he bother to first prove quantum mechanics violates local causality two sections before without using any inequality or hidden variables? We guess because he agreed with Stapp about the logical loophole of concluding quantum nonlocality directly from his inequality [Stapp 2012]:
@@
Above, “properties alien to the original system” rigorously mean variables that do not legitimately pertain to quantum mechanics. Although some have observed that the hidden variables can include the quantum state [Norsen 2011, Gisin 2012, Laudisa 2018], the problem persists with the other “additional variables.” As we observe in the section 3.1.3, the Bell inequality cannot be formulated without additional variables foreign to quantum mechanics, notwithstanding that one of those variables may include the quantum state as indicated by Bell himself [Bell 1981].
2.2.3 Bertlmann’s socks
In 1981 Bell wrote his celebrated paper “Bertlmann’s socks and the nature of reality” [Bell 1981]. On this occasion, Bell did not explicitly prove quantum nonlocality before formulating the inequality. He based his arguments on EPR, but he also mentioned common causes to explain local correlations.
This is one of the papers where he left ambiguous whether his inequality violation should be interpreted as proof of quantum nonlocality. Can we assume that Bell changed his mind about the meaning of his inequality? We do not think so because, in his last paper (cf. 2.2.4), he returned to his previous formulations, i.e., either assuming [Bell 1964] or proving [Bell, Shimony, Horne, and Clauser 1985] quantum nonlocality without introducing hidden variables or mentioning any inequality.
In “Bertlmann’ socks”, Bell chose intuition and ease of interpretation over logical rigor. In Bell’s own words, this paper was one of those that [Bell and Aspect 2004]:
@@ 2.2.3 Bertlmann’s socks
That is why he spent great effort explaining the difference between quantum and classical entanglement through naive analogies, such as those of the socks of Mr. Bertlmann. It is rather surprising that some scientists still believe that quantum entanglement admits such a trivial explanation (cf. sec. 5.2.1).
2.2.4 La nouvelle cuisine
This is Bell’s last paper which appeared in 1990 [Bell 2001]. Here again, Bell’s view of his inequality and quantum nonlocality is crystal clear.
This time his approach differs from his 1975 formulation in “The theory of local beables” in that he proves that quantum mechanics violates both nonlocality criteria, EPR and also local causality. However, what remains unchanged is that he proved quantum nonlocality in section 8 before formulating his inequality.
@@ 2.2.4 La nouvelle cuisine
Although some interpret the above expression as proving that nature itself, including quantum mechanics, is nonlocal, the order in which he presents his argument is unambiguous. First, establishing quantum nonlocality without any inequality, and then, in a separate section, inferring his conclusion through his inequality.
@@
2.2.5 Further writings
In 1975, an editor asked Bell to write a response letter [Bell 1975]. He started the paper by expressing:
@@ 2.2.5 Further writings
In 1977 Bell wrote Free variables and local causality [Bell 1977]. He started the paper by writing:
@@
Once more, his expression openly shows his reasoning. He considered the statement “quantum mechanics is not locally causal” to be before, different from, and not a consequence of, the assertion “cannot be embedded in a locally causal theory” which was the conclusion he used to draw from the violation of his inequality as an additional characteristic of quantum mechanics besides nonlocality, as he explicitly did it in Refs. [Bell, Shimony, Horne, and Clauser 1985, Bell 2001] and we explain in sections 2.2.2 and 2.2.4.
3 Bell’s Proof of Quantum Nonlocality
In this section, we analyze Bell’s explicit argument on the nonlocal character of quantum mechanics. In his 1975 paper, “The theory of local beables” [Bell, Shimony, Horne, and Clauser 1985], Bell gave an explicit argument for quantum nonlocality for the first time. This paper has four outstanding characteristics that were missing in 1964:
@@ 3 Bell’s Proof of Quantum Nonlocality
3.1 Local Causality and Quantum Nonlocality
Firstly, we briefly review the meaning of LC. Unfortunately, as observed by Norsen [Norsen 2011], LC is a little-known concept, and we would add that, if known, it is often misunderstood. For this reason, appendix 5.1 contains a more detailed explanation for those unfamiliar with it. Secondly, like Bell, we prove that quantum mechanics violates LC without the Bell inequality. Thirdly, we analyze the origin of the argumentative conflict between localists and nonlocalists.
3.1.1 Local causality
@@
(2) means that after all factors (known and unknown) are included, whatever Bob decides to do in his distant laboratory cannot influence Alice’s local measurements; and vice versa. Therefore, (2) must be satisfied by locally causal theories. Note that $\lambda$ represents local common causes in general, not necessarily “preexisting values” and can well include the quantum state (cf. appendix 5.3).
3.1.2 Quantum nonlocality